# Flow is slow at the nanoscale: Revisiting the Green–Kubo relation for friction

### v1

@@ Abstract
A central aim of statistical mechanics is to establish connections between a system’s microscopic fluctuations and its macroscopic response to a perturbation. For non-equilibrium transport properties, this amounts to establishing Green–Kubo (GK) relationships. In hydrodynamics, relating such GK expressions for liquid–solid friction to macroscopic slip boundary conditions has remained a long-standing problem due to two challenges: (i) The GK running integral of the force autocorrelation function decays to zero rather than reaching a well-defined plateau value; and (ii) debates persist on whether such a transport coefficient measures an intrinsic interfacial friction or an effective friction in the system. Inspired by ideas from the coarse-graining community, we derive a GK relation for liquid–solid friction where the force autocorrelation is sampled with a constraint of momentum conservation in the liquid. Our expression does not suffer from the “plateau problem” and unambiguously measures an effective friction coefficient, in an analogous manner to Stokes’ law. We further establish a link between the derived friction coefficient and the hydrodynamic slip length, enabling a straightforward assessment of continuum hydrodynamics across length scales. We find that continuum hydrodynamics describes the simulation results quantitatively for confinement length all the way down to 1 nm. Our results also make clear that water flow under nano-confinement is orders of magnitude slower compared to the macroscopic case. Our approach amounts to a straightforward modification to the present standard method of quantifying interfacial friction from molecular simulations, making possible a sensible comparison between surfaces of vastly different slippage.
@@ I Introduction
From a macroscopic perspective, such a property is the slip length $b$, measuring the distance beyond the surface of the wall at which the fluid velocity $v_{x}(z)$ extrapolates to zero. The slip length enters a continuum hydrodynamics description, hereafter referred to as “classical hydrodynamic theory” (CHT), as a boundary condition to the Navier–Stokes equations, which for an interface lying on the $z=z_{0}$ plane, reads
@@
relating the velocity gradient with the slip velocity $v_{\mathrm{s}}=v_{x}(z_{0})$. The no-slip boundary condition $b=0$ applies for surfaces that are highly sticky while for atomically smooth and non-wetting surfaces where there is finite slippage at the interface, $b>0$. The gradient of the fluid velocity is linearly related to the viscous stress of the fluid $\sigma_{xz}=-\eta(\partial v_{x}/\partial z)$, where $\eta$ is the shear viscosity. The stress is itself balanced by the friction force from the solid to the liquid $F_{x}$ per unit surface area $\mathcal{A}$. This allows a linear relation between $F_{x}$ and $v_{\mathrm{s}}$ to be written
@@
From a microscopic perspective, frictional effects coming from dissipation in the liquid manifest through balance of forces at non-equilibrium steady states. A natural question to ask is if one can relate these transport coefficients to equilibrium fluctuations in the liquid, invoking Onsager’s regression hypothesis. In this context, one attempts to seek Green–Kubo (GK) expressions for a liquid–solid friction coefficient, allowing the hydrodynamic boundary conditions to be characterized from a single equilibrium molecular dynamics (EMD) simulation. The seminal work by Bocquet and Barrat[25, 26] (BB) proposed such a relation for $\lambda_{\rm intr}$ given as
@@
However, two major challenges arise upon application of Eq. 4 to measure friction at the liquid–solid interface. The first issue is the well-documented “plateau problem,” [27, 28, 29, 30, 31, 32, 33, 34] referring to the fact that the integral in Eq. 4 in general will decay to zero at long time. While this behavior is often attributed to the swapping of the order in which the thermodynamic and long-time limits are taken [29, 33, 35, 30, 26], we will show—as becomes clear when employing a Mori–Zwanzig formalism to derive the appropriate Langevin dynamics—that this is in fact a direct consequence of using real forces rather than projected forces in Eq. 4. The plateau problem is most severe when the surface is strongly wetting or when the liquid film is very thin (see the SI).
@@ II.1 Mori–Zwanzig projection operator formalism
$\frac{\mathrm{d}A(t)}{\mathrm{d}t}=i\varOmega A(t)-\int^{t}_{0}\!\mathrm{d}t^{\prime}\,K(t-t^{\prime})A(t)+F^{\mathrm{R}}(t),$ (9)
@@ II.2 Applying the operator formalism to liquid–solid friction
Figure 1: The Green–Kubo friction integral. For a liquid–solid interface with high wettability (see Sec. IIIB), we show (a) the force–momentum correlation function, (b) force–force autocorrelation function and (c) the friction GK integral. Using real dynamics results in a plateau problem since the force is correlated with the momentum, leading to negative correlation in the force autocorrelation; this causes the GK integral to vanish at long times. Constrained dynamics solves this issue by enabling the calculation of projected forces for friction, ensuring that the integral plateaus at a finite value
To describe liquid–solid friction, we will consider the problem of a liquid droplet’s stochastic motion on a solid surface, schematically shown in Fig. 1(a). We will work in the solid’s frame of reference so that we can consider it to be at rest. The liquid droplet comprises $N$ particles, whose positions and momenta are $\{\bm{\mathit{r}}_{i},\bm{\mathit{p}}_{i}\}$ and has a total mass of $M=\sum^{N}_{i=1}m_{i}$. For simplicity, we will focus only on the droplet’s motion in the $x$ direction, with a center-of-mass velocity of $\overline{v}_{x}$.
@@
where $\bm{\mathit{f}}_{i}$ is the force on each liquid particle. We will refer to the dynamics associated with $i\mathcal{L}$ as the “real dynamics,” i.e., those that result from propagating Newton’s equations of motions in standard MD simulations. Upon application of the result in Eq. 13, which implies a Markovian approximation, the liquid motion can be described with
@@
According to Eq. 19, to obtain $\lambda_{\rm eff}$, the force autocorrelation function should be sampled with projected dynamics propagated by $i\mathcal{QL}$. However, such a dynamical scheme is not readily realizable in MD simulations, since in general it is not possible write the projection operator $\mathcal{Q}$ explicitly. To make progress, a common though ad-hoc assumption to make is that the random force evolves with $i\mathcal{L}$ instead of $i\mathcal{QL}$, i.e $F^{\mathrm{R}}(t)\approx\mathrm{e}^{i\mathcal{L}t}F_{x}(0)$. The BB formula given by Eq. 4 amounts to such an approximation. Though the GK formula in Eq. 4 can be sampled directly in a MD simulation, a consequence of replacing $i\mathcal{Q}\mathcal{L}$ with $i\mathcal{L}$, is that the random force is no longer uncorrelated with the coarse-grained variable $P_{x}$, violating the condition stated in Eq. 12. The effect is easily seen in the results of MD simulations. For water film $\sim 2.7\,$nm in thickness in contact with a strongly wetting surface (see Sec. III.2), we see in Fig. 1(b) that $\langle F_{x}(\tau)P_{x}(0)\rangle$ initially shows a large positive correlation, exhibiting a weak negative correlation at longer times. Such correlation between the force and the momentum leads to negative contribution to the force autocorrelation function, as shown in Fig. 1(c), making the integral in Eq. 4 vanish at long times, as seen in Fig. 1(d). We can also understand why the plateau problem encountered in the use of the BB formula is most severe for surfaces of strong wettability or thin liquid films is very thin. In these cases, since a large portion of momentum in the liquid can be transfered to the solid, $F_{x}$ and $P_{x}$ are strongly correlated in the real dynamics, and Eq. 12 is severely violated. Moreover, the Markovian approximation that relaxation of $P_{x}$ is much slower than all the other degrees of freedom also becomes more severe.
@@ II.3 Momentum conservation constraint to overcome the plateau problem
In fact, the plateau problem has been discussed in the coarse-graining community, where the issues associated with replacing projected dynamics with real dynamics, along with the limitations of the Markovian approximation are more widely recognized [53, 54]. There have also been developments of algorithms for obtaining projected observables of the GLE directly [55, 56, 57]. Here, we adopt the strategy proposed in Ref. 53: instead of modeling the real dynamics directly, where the Markovian approximation is only good in certain limits, we aim to construct a constrained dynamics in which the Markovian approximation is enforced. It then becomes a modeling question of how well the constrained dynamics represents the physical behavior of interest. In the context of liquid–solid friction, we validate our results directly against NEMD simulations.
@@
An important consequence of the constrained dynamics is that since the liquid momentum $P_{x}$ is conserved, all other degrees of freedom of the system become uncorrelated with $P_{x}$. In other words, they all lie on the subspace orthogonal to $P_{x}$ such that the effect of the projection operator $\mathcal{Q}$ is already encapsulated, i.e.
@@
$\begin{split}\lambda_{\mathrm{eff}}&=\frac{\beta}{\mathcal{A}}\int^{\infty}_{0}\!\mathrm{d}t\,(\mathrm{e}^{i\mathcal{L}_{\mathrm{c}}t}F_{x},F_{x})\\ &\equiv\frac{\beta}{\mathcal{A}}\int^{\infty}_{0}\!\mathrm{d}t\,\langle F_{x}(t)F_{x}(0)\rangle_{\mathrm{c}},\end{split}$ (24)
where in the second line, we have used $\langle\ldots\rangle_{\rm c}$ to denote a canonical average in the constrained system. The lack of correlation in $F_{x}$ and $P_{x}$ under constrained dynamics, seen in Fig. 1(b), means that Eq. 12 is always satisfied and there is no negative contribution to the force autocorrelation function, as seen in Fig. 1(c). Therefore, the integral in Eq. 24 always reaches a well-defined plateau value and does not decay to zero, as seen in Fig. 1(d). We will show in Sec. III.1 that $\lambda_{\rm eff}$ measures an effective friction of the entire system and not, in general, the liquid-solid interfacial friction.
Eq. 24 is a core result of the paper, presenting a GK relation for liquid–solid friction where the force autocorrelation function is sampled with a zero momentum constraint on the liquid. This constraint ensures that the sampled force is already an orthogonal projection of the momentum. While we cannot prove these constrained dynamics are a faithful model for liquid–solid friction, in addition to numerically verifying its usefulness in Sec. III, we can provide several suggestions as to why this might be the case. First, since the constraint does not change the internal energy of the system, the hydrodynamic boundary remains unchanged between real and constrained dynamics, as verified by simulations (see Fig. S3). Second, although constraining the liquid droplet to remain at rest effectively accelerates the microscopic noise contributing to the random force, at the microscopic scale each liquid particle on average experiences the same potential energy surface with the same thermal fluctuations, so the overall dissipation in the system should be largely unaffected provided that liquid droplet is large enough.
@@ III.1 The effective friction coefficient is not an intrinsic property of the interface
The friction coefficient $\lambda_{\rm eff}$ quantifies the dissipation in the system when there is a net relative motion between the the liquid and the solid. When the liquid is pushed out of equilibrium, in general, the velocity of the liquid will not be constant across the cross-section perpendicular to the direction of the flow. Therefore, in addition to dissipation coming from collisions of liquid particles with the solid at the interface, there is also contribution from viscous forces arising between adjacent layers of the fluid flowing at different velocities away from the interface. This means that $\lambda_{\mathrm{eff}}$ will depend on the amount of liquid present, and is not an intrinsic surface property. In fact, the effective friction will only be equal to intrinsic friction $\lambda_{\mathrm{eff}}\approx\lambda_{\mathrm{intr}}$ either in the limit of a very thin film of liquid or when there is perfect slip ($b=\infty$) at the interface and the velocity profile becomes plug-like, such that the fluid velocity is well approximated by the slip velocity $\overline{v}_{x}\approx v_{\mathrm{s}}$. In this limit, Eq. 17 becomes
@@
At this point, it is instructive to point out the difference in our derivation compared to that of BB in Ref. 26. So far, we had worked in the solid’s frame of reference and treated the liquid’s motion as the coarse-grained variable. We then applied the Langevin equation under the Markovian approximation and took the limit where the velocity profile in the fluid is a constant (plug flow) to arrive at Eq. 25. Meanwhile, BB started out by considering the solid wall’s motion as the coarse-grained variable, analogous to extending Brownian motion to a planar geometry. In fact, the same expression for the friction as Eq. 24 would result if we chose to work in the liquid’s frame of reference and had written the Langevin equation
@@
where $M_{\mathrm{w}}$ and $U$ are the mass and velocity of the solid wall, respectively. As discussed above, when the flow profile is plug-like, $U\approx-v_{\mathrm{s}}$ and $\lambda_{\mathrm{eff}}\approx\lambda_{\mathrm{intr}}$. In this limit, we recover the Langevin equation as written by BB in Ref. 26
$M_{\mathrm{w}}\frac{\mathrm{d}U(t)}{\mathrm{d}t}=\lambda_{\mathrm{eff}}\mathcal{A}v_{\mathrm{s}}(t)+F^{\mathrm{R}}(t).$ (27)
@@
Figure 2: Connection between microscopic friction and macroscopic hydrodynamics. (a) Verification of the constitutive relation in Eq. 30 between the total frictional force and the mean fluid velocity for the effective friction. The data points are from NEMD simulations of Poiseuille flow in 2D channels of water confined with surfaces of different wetting coefficients $\alpha$. The lines gives the gradient $1/\lambda_{\mathrm{eff}}$ obtained from the GK relation in EMD simulations. (b) The intrinsic friction coefficients obtained from our approach (GK+CHT) given as $\lambda_{\rm intr}=\eta/b$, with the slip length $b$ obtained from CHT by Eq. 32, which is combined with the GK relation in Eq. 33, yields excellent agreement with reference NEMD simulations for all values of $\alpha$ considered. Meanwhile, the performance of the BB formula, either by taking the long time limit $\lambda_{\mathrm{BB}}(\tau\to\infty)$ or taking the maximum value $\mathrm{max}[\lambda_{\mathrm{BB}}]$, gets worse as the surface becomes more attractive. (c) Very good agreement is also obtained for the predicted velocity profile using our approach (GK+CHT) and reference NEMD data, shown for the case $\alpha=3$ for both Poiseuille (left) and Couette (right) flow.
@@
where $\mathcal{H}_{\mathrm{c}}$ is the unperturbed Hamiltonian and $\overline{v}_{x}\hat{\bm{\mathit{e}}}_{x}$ plays the role of the external field that couples to the liquid momentum as the conjugate variable. To make a connection with $\lambda_{\mathrm{eff}}$ in Eq. 24, we consider the response of $F_{x}$ under constrained dynamics. In this case, the momentum conservation constraint would maintain the liquid’s linear momentum at zero $P_{x}=0$ at equilibrium and at $P_{x}=M\overline{v}_{x}$ out-of-equilibrium. Using linear response theory[58] (detailed in the SI), we show that the response of the tangential frictional force to a small but finite perturbation $\overline{v}_{x}\hat{\bm{\mathit{e}}}_{x}$ is
@@
To verify Eq. 30, we considered systems with water confined in four symmetric channels made up of solid substrates with attractive strengths of $\varepsilon_{\rm wf}=\alpha\varepsilon_{0}$ where the “wetting coefficients” are $\alpha=1,2,3,4$ and $\varepsilon_{0}=1.57\,\mathrm{kJ\,mol^{-1}}$. We also consider one asymmetric channel with $\alpha=1$ for the top wall and $\alpha=4$ for the bottom wall. Various channel heights are considered $H/\mathrm{nm}\approx 1.4,2.7,5.2$ between the first atomic planes of the solids. Details on the systems and molecular models can be found in the SI. Using constrained EMD simulations, we computed $\lambda_{\mathrm{eff}}$ from the GK relation in Eq. 24. To generate a finite relative average velocity of the fluid relative to the solid, we performed NEMD simulations in which a body force is applied on each oxygen atom to mimic the effect of a pressure gradient, generating a Poiseuille flow. The agreement between $F_{x}/\mathcal{A}$ as a function of $\overline{v}_{x}$ from NEMD and $\lambda_{\mathrm{eff}}$ from constrained EMD simulations is robust in the linear response regime across all systems considered, shown for channels of different attractive strengths Fig. 2(a), and for channels of different heights in the SI. In the SI, we also confirmed the equivalence between the equilibrium and out-of-equilibrium force correlation, as expected from the fluctuation–dissipation theorem in the linear response regime.

### v2

# Revisiting the Green–Kubo relation for friction in nanofluidics

@@ Abstract
A central aim of statistical mechanics is to establish connections between a system’s microscopic fluctuations and its macroscopic response to a perturbation. For non-equilibrium transport properties, this amounts to establishing Green–Kubo (GK) relationships. In hydrodynamics, relating such GK expressions for liquid–solid friction to macroscopic slip boundary conditions has remained a long-standing problem due to two challenges: (i) The GK running integral of the force autocorrelation function decays to zero rather than reaching a well-defined plateau value; and (ii) debates persist on whether such a transport coefficient measures an intrinsic interfacial friction or an effective friction in the system. Inspired by ideas from the coarse-graining community, we derive a GK relation for liquid–solid friction where the force autocorrelation is sampled with a constraint of momentum conservation in the liquid. Our expression does not suffer from the “plateau problem” and unambiguously measures an effective friction coefficient, in an analogous manner to Stokes’ law. We further establish a link between the derived friction coefficient and the hydrodynamic slip length, enabling a straightforward assessment of continuum hydrodynamics across length scales. We find that continuum hydrodynamics describes the simulation results quantitatively for confinement length scales all the way down to 1 nm. Our approach amounts to a straightforward modification to the present standard method of quantifying interfacial friction from molecular simulations, making possible a sensible comparison between surfaces of vastly different slippage.
@@ I Introduction
From a macroscopic perspective, such a property is the slip length $b$, measuring the distance beyond the surface of the wall at which the fluid velocity $v_{x}(z)$ extrapolates to zero. The slip length enters a continuum hydrodynamics description, hereafter referred to as “classical hydrodynamic theory” (CHT), as a boundary condition to the Navier–Stokes equations
@@
relating the velocity gradient with the slip velocity $v_{\mathrm{s}}=v_{x}(z_{0})$. The $z=z_{0}$ plane is where the hydrodynamic boundary position is placed. The no-slip boundary condition $b=0$ applies for surfaces that are highly sticky, while for atomically smooth and non-wetting surfaces where there is finite slippage at the interface, $b>0$. The gradient of the fluid velocity is linearly related to the viscous stress of the fluid $\sigma_{xz}=-\eta(\partial v_{x}/\partial z)$, where $\eta$ is the shear viscosity (which we assume takes its bulk value in all regions occupied by the fluid). The stress is itself balanced by the friction force from the solid to the liquid $F_{x}$ per unit surface area $\mathcal{A}$. This allows a linear relation between $F_{x}$ and $v_{\mathrm{s}}$ to be written
@@
From a microscopic perspective, frictional effects coming from dissipation in the liquid manifest through balance of forces at non-equilibrium steady states. A natural question to ask is if one can relate these transport coefficients to equilibrium fluctuations in the liquid, invoking Onsager’s regression hypothesis. In this context, one attempts to seek Green–Kubo (GK) expressions for a liquid–solid friction coefficient, allowing the hydrodynamic boundary conditions to be characterized from a single equilibrium molecular dynamics (EMD) simulation. The seminal work by Bocquet and Barrat[25, 26] (BB) proposed such a relation for $\lambda_{\rm intr}$, given as
@@
However, two major challenges arise upon application of Eq. 4 to measure friction at the liquid–solid interface. The first issue is the well-documented “plateau problem,” [27, 28, 29, 30, 31, 32, 33, 34] referring to the fact that the integral in Eq. 4 in general will decay to zero at long time. While this behavior is often attributed to the swapping of the order in which the thermodynamic and long-time limits are taken [29, 33, 35, 30, 26], we will show—as becomes clear when employing a Mori–Zwanzig formalism to derive the appropriate Langevin dynamics—that this is in fact a direct consequence of using real forces rather than projected forces in Eq. 4. The plateau problem is most severe when the surface is strongly wetting or when the liquid film is very thin (see Fig. S2 of the supplementary material).
@@ II.1 Mori–Zwanzig projection operator formalism
$\frac{\mathrm{d}A(t)}{\mathrm{d}t}=i\varOmega A(t)-\int^{t}_{0}\!\mathrm{d}t^{\prime}\,K(t-t^{\prime})A({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t^{\prime}})+F^{\mathrm{R}}(t),$ (9)
@@ II.2 Applying the operator formalism to liquid–solid friction
Figure 1: The Green–Kubo friction integral. (a) Schematic of the system set up that we use for the derivation; a liquid droplet of total mass $M$ moves stochastically on a solid surface. Its center-of-mass velocity along one direction is indicated by $\bar{v}_{x}$, and the frictional force with the surface is $F_{x}$. For a liquid–solid interface with high wettability (see Sec. III.2), we show (b) the force–momentum correlation function, (c) force–force autocorrelation function and (d) the friction GK integral. Using real dynamics results in a plateau problem since the force is correlated with the momentum, leading to negative correlation in the force autocorrelation; this causes the GK integral to vanish at long times. Constrained dynamics bypasses this issue by imitating the effect of projected forces, ensuring that the integral plateaus to a finite value.
To describe liquid–solid friction, we will consider the problem of a liquid droplet’s stochastic motion on a solid surface, schematically shown in Fig. 1(a). We choose this arrangement both for conceptual simplicity, and to make the connection between this derivation and that for Brownian motion clear; the resulting framework, however, is readily applied to systems comprising a fluid confined between walls. We will work in the solid’s frame of reference so that we can consider it to be at rest. The liquid droplet comprises $N$ particles, whose positions and momenta are $\{\bm{\mathit{r}}_{i},\bm{\mathit{p}}_{i}\}$ and has a total mass of $M=\sum^{N}_{i=1}m_{i}$. For simplicity, we will focus only on the droplet’s motion in the $x$ direction, with a center-of-mass velocity of $\overline{v}_{x}$.
@@
where $\bm{\mathit{f}}_{i}$ is the force on each liquid particle. We will refer to the dynamics associated with $i\mathcal{L}$ as the “real dynamics,” i.e., those that result from propagating Newton’s equations of motion in standard MD simulations. Upon application of the result in Eq. 13, which implies a Markovian approximation, the liquid motion can be described with
@@
According to Eq. 19, to obtain $\lambda_{\rm eff}$, the force autocorrelation function should be sampled with projected dynamics propagated by $i\mathcal{QL}$. However, such a dynamical scheme is not readily realizable in MD simulations, since in general it is not possible write the projection operator $\mathcal{Q}$ explicitly. To make progress, a common though ad-hoc assumption to make is that the random force evolves with $i\mathcal{L}$ instead of $i\mathcal{QL}$, i.e $F^{\mathrm{R}}(t)\approx\mathrm{e}^{i\mathcal{L}t}F_{x}(0)$. The BB formula given by Eq. 4 amounts to such an approximation. Though the GK formula in Eq. 4 can be sampled directly in a MD simulation, a consequence of replacing $i\mathcal{Q}\mathcal{L}$ with $i\mathcal{L}$, is that the random force is no longer uncorrelated with the coarse-grained variable $P_{x}$, violating the condition stated in Eq. 12. The effect is easily seen in the results of MD simulations. For a water film $\sim 2.7\,$nm in thickness in contact with a strongly wetting surface (see Sec. III.2), we see in Fig. 1(b) that $\langle F_{x}(\tau)P_{x}(0)\rangle$ initially shows a large positive correlation, exhibiting a weak negative correlation at longer times. Such correlation between the force and the momentum leads to a negative contribution to the force autocorrelation function, as shown in Fig. 1(c), making the integral in Eq. 4 vanish at long times, as seen in Fig. 1(d). We can also understand why the plateau problem encountered in the use of the BB formula is most severe for surfaces of strong wettability or when the liquid film is very thin. In these cases, since a large portion of momentum in the liquid can be transfered to the solid, $F_{x}$ and $P_{x}$ are strongly correlated in the real dynamics, and Eq. 12 is severely violated. Moreover, the Markovian approximation that relaxation of $P_{x}$ is much slower than all the other degrees of freedom also becomes more severe.
@@ II.3 Momentum conservation constraint to overcome the plateau problem
In fact, the plateau problem has been discussed in the coarse-graining community, where the issues associated with replacing projected dynamics with real dynamics, along with the limitations of the Markovian approximation, are more widely recognized [53, 54]. There have also been developments of algorithms for obtaining projected observables of the GLE directly [55, 56, 57]. Here, we adopt the strategy proposed in Ref. 53: instead of modeling the real dynamics directly, where the Markovian approximation is only good in certain limits, we aim to construct a constrained dynamics in which the Markovian approximation is enforced. It then becomes a modeling question of how well the constrained dynamics represents the physical behavior of interest. In the context of liquid–solid friction, we validate our results directly against NEMD simulations.
@@
An important consequence of the constrained dynamics is that since the liquid momentum $P_{x}$ is conserved, all other degrees of freedom of the system become uncorrelated with $P_{x}$. In other words, they all lie on the subspace orthogonal to $P_{x}$ such that the effect of the projection operator $\mathcal{Q}$ is already encapsulated, i.e.,
@@
$\begin{split}\lambda_{\mathrm{eff}}&=\frac{\beta}{\mathcal{A}}\int^{\infty}_{0}\!\mathrm{d}t\,(\mathrm{e}^{i\mathcal{L}_{\mathrm{c}}t}F_{x},F_{x})\\ &\equiv\frac{\beta}{\mathcal{A}}\int^{\infty}_{0}\!\mathrm{d}t\,\langle F_{x}(t)F_{x}(0)\rangle_{\mathrm{c}}.\end{split}$ (24)
In the second line, we have used $\langle\ldots\rangle_{\rm c}$ to denote a canonical average in the constrained system. The lack of correlation in $F_{x}$ and $P_{x}$ under constrained dynamics, seen in Fig. 1(b), means that Eq. 12 is always satisfied and there is no negative contribution to the force autocorrelation function, as seen in Fig. 1(c). Therefore, the integral in Eq. 24 always reaches a well-defined plateau value and does not decay to zero, as seen in Fig. 1(d). We will show in Sec. III.1 that $\lambda_{\rm eff}$ measures an effective friction of the entire system and not, in general, the liquid-solid interfacial friction.
Eq. 24 is a core result of the paper, presenting a GK relation for liquid–solid friction where the force autocorrelation function is sampled with a zero momentum constraint on the liquid. We stress that the constraint is proposed as a simple way to ensure that the sampled force obeys the orthogonality condition (Eq. 12), such that the friction coefficient can be obtained from EMD, without explicitly calculating projected forces. While this constrained ensemble is by no means a “physical” model of the system’s dynamics, in addition to numerically verifying its usefulness (see Sec. III), we can provide suggestions as to why it might retain many physical properties relevant for describing liquid–solid friction. First, since the constraint does not change the internal energy of the system, the hydrodynamic boundary and surface roughness[58, 14] remain unchanged between real and constrained dynamics, as verified by simulations (see Fig. S3). Second, when the liquid droplet is constrained to remain at rest, its center-of-mass motion becomes “infinitely” slower compared to the fast collisions individual liquid particles undergo. While this effectively accelerates the microscopic noise contributing to the random force, at the microscopic scale, each liquid particle on average experiences the same potential energy surface with the same thermal fluctuations, so the overall dissipation in the system should be largely unaffected provided that liquid droplet is sufficiently big.
@@ III.1 The effective friction coefficient is not an intrinsic property of the interface
The friction coefficient $\lambda_{\rm eff}$ quantifies the dissipation in the system when there is a net relative motion between the the liquid and the solid. When the liquid is pushed out of equilibrium, in general, the velocity of the liquid will not be constant across the cross-section perpendicular to the direction of the flow. Therefore, in addition to dissipation coming from collisions of liquid particles with the solid at the interface, there are also contributions from viscous forces arising between adjacent layers of the fluid flowing at different velocities away from the interface. This means that $\lambda_{\mathrm{eff}}$ will depend on the amount of liquid present, and is not an intrinsic surface property. In fact, the effective friction will only be equal to the intrinsic friction $\lambda_{\mathrm{eff}}\approx\lambda_{\mathrm{intr}}$ either in the limit of a very thin film of liquid or when there is perfect slip ($b=\infty$) at the interface and the velocity profile becomes plug-like; in such cases, the slip velocity is well approximated by the average fluid velocity, $\overline{v}_{x}\approx v_{\mathrm{s}}$. In this limit, Eq. 17 becomes
@@
At this point, it is instructive to point out the difference in our derivation compared to that of BB in Ref. 26. So far, we have worked in the solid’s frame of reference and treated the liquid’s motion as the coarse-grained variable. We have then applied the Langevin equation under the Markovian approximation and taken the limit where the velocity profile in the fluid is a constant (plug flow) to arrive at Eq. 25. Meanwhile, BB started out by considering the solid wall’s motion as the coarse-grained variable, analogous to extending Brownian motion to a planar geometry. In fact, the same expression for the friction as Eq. 24 would result had we chosen to work in the liquid’s frame of reference and written the Langevin equation as
@@
where $M_{\mathrm{w}}$ and $U$ are the mass and velocity of the solid wall, respectively. When the flow profile is plug-like, $U\to-\overline{v}_{x}\approx-v_{\mathrm{s}}$ and, as discussed above, $\lambda_{\mathrm{eff}}\approx\lambda_{\mathrm{intr}}$. In this limit, we recover the Langevin equation as written by BB in Ref. 26.
@@
Figure 2: Connection between microscopic friction and macroscopic hydrodynamics. (a) Verification of the constitutive relation in Eq. 29 between the total frictional force and the mean fluid velocity for the effective friction. The data points are from NEMD simulations of Poiseuille flow in 2D channels of water confined with surfaces of different wetting coefficients $\alpha$. The solid lines have slopes $1/\lambda_{\mathrm{eff}}$, where $\lambda_{\mathrm{eff}}$ has been obtained from the GK relation with constrained EMD simulations (Eq. 32, label GK*). (b) The slip lengths obtained from Eq. 31, using $\lambda_{\mathrm{eff}}$ from EMD simulations (Eq. 32), labeled GK*+CHT, are in excellent agreement with those using $\lambda_{\mathrm{eff}}$ from linear fits to NEMD data points in (a). Meanwhile, the performance of the BB formula, either by taking the long time limit $\lambda_{\mathrm{BB}}(\tau\to\infty)$ or taking the maximum value $\mathrm{max}[\lambda_{\mathrm{BB}}]$, and interpreting it as an intrinsic friction coefficient (Eq. 3) gets worse as the surface becomes more attractive. (c) Very good agreement is also obtained for the predicted velocity profile using our approach (GK*+CHT) and reference NEMD data, shown for the case $\alpha=3$ for both Poiseuille (left) and Couette (right) flow.
@@
where $\mathcal{H}_{\mathrm{c}}$ is the unperturbed Hamiltonian and $\overline{v}_{x}\hat{\bm{\mathit{e}}}_{x}$ plays the role of the external field that couples to the liquid momentum as the conjugate variable. To make a connection with $\lambda_{\mathrm{eff}}$ in Eq. 24, we consider the response of $F_{x}$ under constrained dynamics. In this case, the momentum conservation constraint would maintain the liquid’s linear momentum at zero $P_{x}=0$ at equilibrium and at $P_{x}=M\overline{v}_{x}$ out-of-equilibrium. Using linear response theory[59] (detailed in the supplementary material), we show that the response of the tangential frictional force to a small but finite perturbation $\overline{v}_{x}\hat{\bm{\mathit{e}}}_{x}$ is
@@
Applying the framework that we have derived to confined fluids is straightforward; the total frictional force on the liquid ($F_{x}$) now contains contributions from all solid surfaces in contact with the fluid, while $v_{x}$ is simply the average flow velocity of the fluid. To verify Eq. 29, we considered systems with water confined in four symmetric channels made up of solid substrates with attractive strengths of $\varepsilon_{\rm wf}=\alpha\varepsilon_{0}$ where the “wetting coefficients” are $\alpha=1,2,3,4$ and $\varepsilon_{0}=1.57\,\mathrm{kJ\,mol^{-1}}$. We also consider one asymmetric channel with $\alpha=1$ for the top wall and $\alpha=4$ for the bottom wall. Various channel heights between the first atomic planes of the solids are considered, $H/\mathrm{nm}\approx 1.4,2.7,5.2$. Details on the systems and molecular models can be found in the supplementary material. Using constrained EMD simulations, we computed $\lambda_{\mathrm{eff}}$ from the GK relation in Eq. 24. For reference NEMD simulations, we performed simulations of Poiseuille flow by applying a body force on each oxygen atom to mimic the effect of a pressure gradient. In the regime of low driving force, the linear response predicted by $\lambda_{\mathrm{eff}}$ obtained from constrained EMD simulations (labeled GK*) is in excellent agreement with $\overline{v}_{x}$ measured from NEMD simulations, shown for channels of different attractive strengths in Fig. 2(a), and for channels of different heights in the supplementary material. In the supplementary material (Fig. S3), we also confirmed the equivalence between the equilibrium and out-of-equilibrium force autocorrelation, as expected from the fluctuation–dissipation theorem in the linear response regime.